For x and y in R, let d(x,y)=(|x-y|)/(1+|x-y|),prove that d defines a.bounded metric on R


1
Expert's answer
2021-07-26T14:22:45-0400

1)

Positivity: "d(x,y)\ge0" with equality if x=y


2)

Symmetry: "d(x,y)=d(y,x)"

"\frac{|x-y|}{1+|x-y|}=\frac{|y-x|}{1+|y-x|}"


3)

Triangle Inequality: "d(x,y)\le d(x,z)+d(z,y)"

"d(x,y)=\frac{|x-y|}{1+|x-y|}=\frac{|(x-z)+(z-y)|}{1+|(x-z)+(z-y)|}\le \frac{|(x-z)+(z-y)|}{1+|(x-z)+(z-y)|}=d(x,z)+d(z,y)"